# Active learning emulators for nuclear two-body scattering in momentum space

### v1

@@ II.3 Greedy snapshot selection
How are the locations of the snapshot calculations determined? Here, we follow the greedy snapshot selection process developed in Ref. [Maldonado:2025ftg] and benchmark it using the proper orthogonal decomposition (POD). We refer the reader to Section IV in Ref. [Maldonado:2025ftg] for a comprehensive discussion of these two complementary approaches to snapshot selection.
@@
Motivated by the findings in Ref. [Maldonado:2025ftg], we assume that $\norm{\mathbf{r}({\bf\it\theta})}$ and $\norm{\mathbf{e}({\bf\it\theta})}$ are approximately proportional to each other, so the factor of proportionality can be determined at the expense of a single additional high-fidelity calculations at the end of the greedy iteration, as discussed in Section IV.A.3 of Ref. [Maldonado:2025ftg]. Under this assumption, the error estimator can hence be calibrated without performing expensive FOM calculations during the emulator’s online stage. A rigorous upper bound of the emulator error, which would constitute a conservative error estimate, was derived in Equation (48) in Ref. [Maldonado:2025ftg]. However, evaluating it requires a computationally expensive estimate of the minimum singular value of $\mathbf{A}({\bf\it\theta})$, e.g., using the successive constraint method (SCM) [HUYNH2007473].
@@ III.1 Test case: Minnesota potential
As a test case, we build an emulator for the Minnesota potential [THOMPSON197753] in the $^{1}\text{S}_{0}$ channel, which, in coordinate space, is given by,
@@
where $r$ is the relative coordinate. It was fitted in Ref. [THOMPSON197753] to match experimentally extracted effective range parameters, leading to the following best-fit model parameters in Eq. (21): $\theta_{1}\equiv V_{R}=200\,\text{MeV}$, $\theta_{2}\equiv V_{s}=-91.85\,\text{MeV}$, $\kappa_{R}=1.487~\text{fm}^{-2}$, and $\kappa_{s}=0.465~\text{fm}^{-2}$. In the following, we will vary the two affine parameters $(V_{R},V_{s})$ while keeping the non-affine parameters $(\kappa_{R},\kappa_{s})$ fixed at their best-fit values, and $\theta_{0}\equiv 0$. Because our FOM and ROM solvers are based on the LS equation, we transform Eq. (21) to momentum space by carrying out the Fourier transform analytically, resulting in: [^2] We follow the convention $\Braket{\mathbf{r}|\mathbf{p}}=e^{i\mathbf{r}\cdot\mathbf{p}}/(2\pi)^{3/2}$.
@@
Figure 1: Convergence of the greedy algorithm for the Minnesota potential (22) at $E_{\text{lab}}=20\,\text{MeV}$ in the $^{1}\text{S}_{0}$ channel. Each panel shows a color map of the true error (17) in the two-dimensional parameter space of the potential (see the legends). The pound symbol “#” specifies the number of greedy iterations passed: Panel (a) depicts the initial configuration, for which we randomly select snapshots in the parameter space using LHS. The remaining panels (b) through (f) show the configuration in the emulator after 2, 4, 6, 8, and 9 iterations, respectively. Note that not all greedy iterations are depicted. As shown, the greedy algorithm places additional snapshots at locations where the (estimated) emulator error is largest, thereby iteratively reducing the emulator error locally and globally. Note that the emulator error vanishes at the origin (which is not a grid point in the panels) because the interaction is zero as there is no parameter-independent term, rendering the ROM equations homogeneous. See the main text for details.
@@
Figure 1 shows the greedy algorithm in action for the Minnesota potential (22) at a representative laboratory energy, $E_{\text{lab}}=20\,\text{MeV}$. [^3] See also Figure 6 in Ref. [Maldonado:2025ftg] for a similar, albeit one-dimensional, illustration of the greedy algorithm in action. Each panel in Fig. 1 depicts a color map of the true error (17) in the two-dimensional parameter space of the potential (see the legends). Specifically, Fig. 1a) shows the initial configuration, for which we randomly select two snapshots in the parameter space using LHS. The black squares near the center of the shown parameter space highlight the snapshot location. At these points, the emulator error vanishes, as expected.
@@
From this initial configuration, we let the greedy algorithm improve the snapshot basis by adding one snapshot per iteration: Panels b) through f) show the color map of the actual error after 2, 4, 6, 8, and 9 of those iterations, respectively, as indicated by the number following the pound symbol (“#”) in the annotated text. Note that, for brevity, not all iterations are depicted in Fig. 1. At the end of each iteration, we (re)orthonormalize the emulator basis by applying the QR decomposition to the snapshot matrix $\mathbf{X}$, as described in Ref. [Maldonado:2025ftg]. An orthonormal emulator basis significantly improves the numerical stability of the ROM. The greedy algorithm terminates after nine iterations (see panel f) because all errors shown are below the requested threshold, here set to $\varepsilon<10^{-8}$. After these iterations, the emulator basis consists of $2+9=11$ snapshots (corresponding to the dark points in panel f).
@@
As illustrated by Fig. 1, the greedy algorithm enhances a given snapshot basis by placing additional snapshots in locations where the estimated emulator error is the largest. The estimated location of maximum error in this application closely matches the actual location. The greedy algorithm thus effectively reduces the emulator error locally and globally until the requested tolerance $\alpha$ is met. Furthermore, we observe in Fig. 1 that the greedy algorithm adds snapshot calculations at the boundaries of the parameter space to the emulator basis. This behavior was already observed in Ref. [Maldonado:2025ftg]. In these regions, the emulator must extrapolate so that larger errors can be expected. While in two-dimensional parameter spaces one can manually place snapshot locations at the boundaries to improve the emulator’s accuracy, in higher-dimensional spaces this may no longer be possible due to the curse of dimensionality. On the other hand, the greedy algorithm always places snapshots so that the emulator error is optimally reduced locally, including choosing snapshots at the boundaries that are most important for error reduction.
@@ III.2 Chiral potential at N2LO
Next, we apply the greedy emulators to a realistic chiral potential. Following Refs. [Drischler:2021qoy, Maldonado:2025ftg], we choose here one of the local chiral potentials constructed by Gezerlis et al. in Ref. [Gezerlis:2014zia]: the one in the $np$-channel at N2LO and with coordinate space cutoff $R=1.0\,\text{fm}$ and spectral function cutoff $\tilde{\Lambda}=1000\,\text{MeV}$. We will refer to this potential as the “GT+ potential.” These local chiral potentials are commonly used in quantum Monte Carlo calculations of finite nuclei and infinite matter [Piarulli:2019cqu]. However, we note that the GT+ potential is merely a convenient choice and that our FOM and ROMs are broadly applicable to any local or non-local potential in momentum space. The affine parameters of the GT+ potential are the NN contact LECs ${\bf\it\theta}=\{1,C_{S},C_{T},C_{1},C_{2},...,C_{7}\}$, where we added again an auxiliary dimension to incorporate constant contributions; here, all pionic terms are kept at their best-fit values. The authors of Ref. [Gezerlis:2014zia] provide a C++ code that evaluates their potentials in both coordinate and momentum space. We modified this code to output the affine components $\mathcal{V}_{a}^{(\ell,\ell^{\prime})}(k,k^{\prime})$ in Eq. (4) in momentum space. In the following, we consider variations in all nine LECs to benchmark our emulators, although not all of them contribute to all partial-wave channels, as shown in Appendix I of Ref. [Gezerlis:2014zia].
@@ III.2.1 Emulator convergence
We study the convergence of the greedy algorithm and benchmark it against the POD approach to snapshot selection. For emulator training, we use a set of 150 random sampling points obtained via LHS in the region of $\pm 40\%$ of each LEC’s best-fit value [Gezerlis:2014zia]. Although both have access to the same candidate snapshots, the greedy and POD approaches handle them differently during training. As illustrated in Fig. 1 in Ref. [Maldonado:2025ftg], the POD approach computes high-fidelity solutions at all candidate snapshots and then compresses the emulator basis using a truncated SVD. In contrast, the greedy emulator performs high-fidelity calculations only on a small subset of candidate snapshots, namely those for which the emulator error was estimated to be largest during the greedy iteration. The emulator error can be estimated efficiently without requiring high-fidelity calculations, as discussed in Sec. II.3. For more details, see Section IV.B in Ref. [Maldonado:2025ftg]. After training, we validate the emulators using 500 random points obtained via LHS similar to the candidate snapshots.
Figure 2: Emulator convergence and comparison at $E_{\text{lab}}=100\,\text{MeV}$ (left panel) and $200\,\text{MeV}$ (right panel) for the GT+ potential [Gezerlis:2014zia] in the $^{1}\text{S}_{0}$ channel. The greedy algorithm applied to the G-ROM is compared to the POD approach. Both emulators have access to the 150 candidate snapshots in a nine-dimensional parameter space, randomly selected via LHS in the region of $\pm 40\%$ of the best-fit LEC values (obtained in Ref. [Gezerlis:2014zia]) However, the POD emulator uses all of them, while the greedy emulator selects only a small subset. The $x$-axis shows the size of the emulator basis, corresponding to the snapshot calculations performed (greedy algorithm) or the dominant POD modes used (POD approach). To test their accuracy, we use a validation set of 500, similarly obtained via LHS as the candidate snapshots during training. The boxes indicate the range between the first and third quartiles, and the line within the box represents the median. The whiskers show the range from the 5th to the 95th percentiles. See the main text for details.
@@
Figures 2 and 3 show, using so-called box-and-whisker plots, the distributions of the exact relative errors (17) in the uncoupled $^{1}\text{S}_{0}$ and coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channels, respectively, based on the validation points. [^4] See also Figures 7 through 10 in Ref. [Maldonado:2025ftg] for similar convergence plots in the $^{1}\text{S}_{0}$ channel. The left panels correspond to $E_{\text{lab}}=100\,\text{MeV}$, the right panels to $E_{\text{lab}}=200\,\text{MeV}$. We observe similar convergence patterns for the two ROMs and thus focus here on the G-ROM. The filled boxes indicate the range from the first to the third quartiles, and the line within the box represents the median. In addition, the whiskers show the range from the 5th to the 95th percentiles. For the greedy approach, the $x$-axis specifies the number of snapshot calculations performed in total, whereas for the POD approach, it specifies the number of dominant POD modes. In both cases, it specifies the size of the emulator basis.
@@
As in Ref. [Maldonado:2025ftg], we find that the greedy algorithm achieves similar accuracies than the POD, although its errors are systematically slightly larger. This observation is not surprising because the POD approach has access to the most information and uses, by construction, the most important emulator basis vectors in terms of the singular values (associated with the dominant POD modes). However, the POD approach has two disadvantages over the greedy approach: First, it requires far more high-fidelity calculations than the other emulators, which may not be affordable in computationally more demanding applications, such as three-body scattering (see Ref. [Gnech:2025gsy, Gnech:2025lbg]). Second, it does not estimate errors.
@@ III.2.2 Emulation of phase shifts
Figure 4: High-fidelity (or FOM) and emulated (or ROM) phase shifts and mixing angles as a function of the laboratory energy for the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channel based on the GT+ potential. The panels (a) to (c) show the phase shifts in the $^{3}\text{S}_{1}$ (panel a) and in the $^{3}\text{D}_{1}$ channel (panel b), as well as the mixing angle $\varepsilon_{1}$ (panel c). The solid black lines represent the high-fidelity results, while the dots indicate the emulated results at a few selected energies. The panels (d) to (f) show the relative error (23) of the G-ROM and LSPG-ROM, respectively, with respect to the high-fidelity calculations. This comparison was made for two requested error tolerances on the $\tilde{t}-$matrix, $\alpha=10^{-4}$ and $10^{-7}$, as reported in the legend. See the main text for details.
Next, we emulate phase shifts based on the GT+ potential. To this end, we construct emulators on a grid of laboratory energies for a given partial-wave channel. Figure 4 shows our results for the emulated phase shifts and mixing angles at the potential’s best-fit values of the LECs as a function of the laboratory energy in the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channel. [^5] The best-fit values were obtained in Ref. [Gezerlis:2014zia] using a least-squares fit to scattering phase shifts and the $^{1}\text{S}_{0}$ scattering lengths. We choose the G-ROM and the deuteron channel as a coupled-channel case. The results are representative of the LSPG-ROM and other partial-wave channels we have studied, including the coupled and uncoupled P and F partial waves. [^6] Note that the approach in Ref. [Maldonado:2025ftg] was limited to uncoupled channels and local potentials (in coordinate space), restrictions that our emulators relax.
To obtain the emulated phase shifts in Fig. 4, we train the emulator using a pool of 500 candidate snapshots randomly selected in the region of $\pm 40\%$ of the best-fit LECs values obtained in Ref. [Gezerlis:2014zia]. Given a single initial snapshot randomly chosen from this pool, the greedy algorithm estimates the error of each remaining candidate snapshot, performs a single high-fidelity calculation for the one with the largest (estimated) error, and adds the orthonormalized snapshot to the emulator basis. This process is repeated until the requested error tolerance $\alpha$ specified in the legend is met. After training is complete, we run the emulators using the best-fit LEC values from Ref. [Gezerlis:2014zia].
@@
We conclude that the greedy algorithm enables efficient, accurate training of scattering emulators. Note that second-order corrections from applying the Kohn variational principle (KVP) [Furnstahl:2020abp, Garcia:2023slj] are not needed to achieve this level of accuracy.
In contrast to Ref. [Maldonado:2025ftg], we have not encountered spurious singularities known as Kohn (or Schwartz) anomalies that occur when the emulator equations are singular or near-singular [Drischler:2021qoy]. This may suggest that scattering calculations in momentum space are less prone to such anomalies, regardless of the ROM used, likely because the $t$-matrix is complex-valued, as discussed in Ref. [Drischler:2021qoy]. Only the LSPG-ROM has been analytically shown in Ref. [Maldonado:2025ftg] to be less prone to Kohn anomalies.

### v2

@@ II.3 Greedy snapshot selection
How are the locations of the snapshot calculations determined? Here, we follow the greedy snapshot selection process and benchmark it using the proper orthogonal decomposition (POD). We refer the reader to Section IV in Ref. [28] for a comprehensive discussion of these two complementary approaches to snapshot selection.
@@
Motivated by the findings in Ref. [28], we assume that $\norm{\mathbf{r}({\bf\it\theta})}$ and $\norm{\mathbf{e}({\bf\it\theta})}$ are approximately proportional to each other, so the factor of proportionality can be determined at the expense of a single additional high-fidelity calculations at the end of the greedy iteration, as discussed in Section IV.A.3 of Ref. [28]. Under this assumption, the error estimator can hence be calibrated without performing expensive FOM calculations during the emulator’s online stage. A rigorous upper bound of the emulator error, which would constitute a conservative error estimate, was derived in Equation (48) in Ref. [28]. However, evaluating it requires a computationally expensive estimate of the minimum singular value of $\mathbf{A}({\bf\it\theta})$, e.g., using the successive constraint method [23].
@@ III.1 Test case: Minnesota potential
As a test case, we build an emulator for the Minnesota potential [47] in the $^{1}\text{S}_{0}$ channel, which, in coordinate space, is given by,
@@
where $r$ is the relative coordinate. It was fitted in Ref. [47] to match experimentally extracted effective range parameters, leading to the following best-fit model parameters in Eq. (21): $\theta_{1}\equiv V_{R}=200\,\text{MeV}$, $\theta_{2}\equiv V_{s}=-91.85\,\text{MeV}$, $\kappa_{R}=1.487~\text{fm}^{-2}$, and $\kappa_{s}=0.465~\text{fm}^{-2}$. In the following, we will vary the two affine parameters $(V_{R},V_{s})$ while keeping the non-affine parameters $(\kappa_{R},\kappa_{s})$ fixed at their best-fit values, and $\theta_{0}\equiv 0$. Because our FOM and ROM solvers are based on the LS equation, we transform Eq. (21) to momentum space by carrying out the Fourier transform analytically, resulting in: [^3] We follow the convention $\Braket{\mathbf{r}|\mathbf{p}}=e^{i\mathbf{r}\cdot\mathbf{p}}/(2\pi)^{3/2}$.
@@
Figure 1: Convergence of the greedy algorithm for the Minnesota potential (22) at $E_{\text{lab}}=20\,\text{MeV}$ in the $^{1}\text{S}_{0}$ channel. Each panel shows a color map of the true error (17) in the two-dimensional parameter space of the potential (see the legends). The pound symbol “#” specifies the number of greedy iterations passed: Panel (a) depicts the initial configuration, for which we randomly select snapshots in the parameter space using LHS. The remaining panels (b) through (f) show the configuration in the emulator after 2, 4, 6, 8, and 9 iterations, respectively. Note that not all greedy iterations are depicted. As shown, the greedy algorithm places additional snapshots at locations where the (estimated) emulator error is largest, thereby iteratively reducing the emulator error locally and globally. Note that the emulator error vanishes at the origin (which is not a grid point in the panels) because the interaction is zero as there is no parameter-independent term, rendering the ROM equations homogeneous.
@@
Figure 1 shows the greedy algorithm in action for the Minnesota potential (22) at a representative laboratory energy, $E_{\text{lab}}=20\,\text{MeV}$. [^4] See also Figure 6 in Ref. [28] for a similar, albeit one-dimensional, illustration of the greedy algorithm in action. Each panel in Fig. 1 depicts a color map of the true error (17) in the two-dimensional parameter space of the potential (see the legends). Specifically, Fig. 1a) shows the initial configuration, for which we randomly select two snapshots in the parameter space using LHS. The black squares near the center of the shown parameter space highlight the snapshot location. At these points, the emulator error vanishes, as expected.
@@
From this initial configuration, we let the greedy algorithm improve the snapshot basis by adding one snapshot per iteration: Panels b) through f) show the color map of the actual error after 2, 4, 6, 8, and 9 of those iterations, respectively, as indicated by the number following the pound symbol (“#”) in the annotated text. Note that, for brevity, not all iterations are depicted in Fig. 1. At the end of each iteration, we (re)orthonormalize the emulator basis by applying the QR decomposition to the snapshot matrix $\mathbf{X}$. An orthonormal emulator basis significantly improves the numerical stability of the ROM. The greedy algorithm terminates after nine iterations (see panel f) because all errors shown are below the requested threshold, here set to $\varepsilon<10^{-8}$. After these iterations, the emulator basis consists of $2+9=11$ snapshots (corresponding to the dark points in panel f).
@@
As illustrated by Fig. 1, the greedy algorithm enhances a given snapshot basis by placing additional snapshots in locations where the estimated emulator error is the largest. The estimated location of maximum error in this application closely matches the actual location. The greedy algorithm thus effectively reduces the emulator error locally and globally until the requested tolerance $\alpha$ is met. Furthermore, we observe in Fig. 1 that the greedy algorithm adds snapshot calculations at the boundaries of the parameter space to the emulator basis. This behavior was already observed in Ref. [28]. In these regions, the emulator must extrapolate so that larger errors can be expected. While in two-dimensional parameter spaces one can manually place snapshot locations at the boundaries to improve the emulator’s accuracy, in higher-dimensional spaces this may no longer be possible due to the curse of dimensionality. On the other hand, the greedy algorithm always places snapshots so that the emulator error is optimally reduced locally, including choosing snapshots at the boundaries that are most important for error reduction.
@@ III.2 Chiral potential at N2LO
Next, we apply the greedy emulators to a realistic chiral potential. Following Refs. [8, 28], we choose here one of the local chiral potentials constructed by Gezerlis et al. in Ref. [18]: the one in the $np$-channel at N2LO and with coordinate space cutoff $R=1.0\,\text{fm}$ and spectral function cutoff $\tilde{\Lambda}=1000\,\text{MeV}$. We will refer to this potential as the “GT+ potential.” These local chiral potentials are commonly used in quantum Monte Carlo calculations of finite nuclei and infinite matter [38]. However, we note that the GT+ potential is merely a convenient choice and that our FOM and ROMs are broadly applicable to any local or non-local potential in momentum space for which an affine decomposition (4) is possible. The affine parameters of the GT+ potential are the NN contact LECs ${\bf\it\theta}=\{1,C_{S},C_{T},C_{1},C_{2},...,C_{7}\}$, where we added again an auxiliary dimension to incorporate constant contributions; here, all pionic terms are kept at their best-fit values. The authors of Ref. [18] provide a C++ code that evaluates their potentials in both coordinate and momentum space. We modified this code to output the affine components $\mathcal{V}_{a}^{(\ell,\ell^{\prime})}(k,k^{\prime})$ in Eq. (4) in momentum space. In the following, we consider variations in all nine LECs to benchmark our emulators, although not all of them contribute to all partial-wave channels, as shown in Appendix I of Ref. [18].
@@ III.2.1 Emulator convergence
We study the convergence of the greedy algorithm and benchmark it against the POD approach to snapshot selection. For emulator training, we use a set of 150 random sampling points obtained via LHS in the region of $\pm 40\%$ of each LEC’s best-fit value [18]. Although both have access to the same candidate snapshots, the greedy and POD approaches handle them differently during training. As illustrated in Fig. 1 in Ref. [28], the POD approach computes high-fidelity solutions at all candidate snapshots and then compresses the emulator basis using a truncated SVD. In contrast, the greedy emulator performs high-fidelity calculations only on a small subset of candidate snapshots, namely those for which the emulator error was estimated to be largest during the greedy iteration. The emulator error can be estimated efficiently without requiring high-fidelity calculations, as discussed in Sec. II.3. For more details, see Section IV.B in Ref. [28]. After training, we validate the emulators using 500 random points obtained via LHS similar to the candidate snapshots.
Figure 2: Emulator convergence and comparison at $E_{\text{lab}}=100\,\text{MeV}$ (left panel) and $200\,\text{MeV}$ (right panel) for the GT+ potential [18] in the $^{1}\text{S}_{0}$ channel. The greedy algorithm applied to the G-ROM is compared to the POD approach. Both emulators have access to the 150 candidate snapshots in a nine-dimensional parameter space, randomly selected via LHS in the region of $\pm 40\%$ of the best-fit LEC values (obtained in Ref. [18]) However, the POD emulator uses all of them, while the greedy emulator selects only a small subset. The $x$-axis shows the size of the emulator basis, corresponding to the snapshot calculations performed (greedy algorithm) or the dominant POD modes used (POD approach). To test their accuracy, we use a validation set of 500, similarly obtained via LHS as the candidate snapshots during training. The boxes indicate the range between the first and third quartiles, and the line within the box represents the median. The whiskers show the range from the 5th to the 95th percentiles.
@@
Figures 2 and 3 show, using so-called box-and-whisker plots, the distributions of the exact relative errors (17) in the uncoupled $^{1}\text{S}_{0}$ and coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channels, respectively, based on the validation points. [^5] See also Figures 7 through 10 in Ref. [28] for similar convergence plots in the $^{1}\text{S}_{0}$ channel. The left panels correspond to $E_{\text{lab}}=100\,\text{MeV}$, the right panels to $E_{\text{lab}}=200\,\text{MeV}$. We observe similar convergence patterns for the two ROMs and thus focus here on the G-ROM. The filled boxes indicate the range from the first to the third quartiles, and the line within the box represents the median. In addition, the whiskers show the range from the 5th to the 95th percentiles. For the greedy approach, the $x$-axis specifies the number of snapshot calculations performed in total, whereas for the POD approach, it specifies the number of dominant POD modes. In both cases, it specifies the size of the emulator basis.
@@
As in Ref. [28], we find that the greedy algorithm achieves similar accuracies than the POD, although its errors are systematically slightly larger. This observation is not surprising because the POD approach has access to the most information and uses, by construction, the most important emulator basis vectors in terms of the singular values (associated with the dominant POD modes). However, the POD approach has two disadvantages over the greedy approach: First, it requires far more high-fidelity calculations than the other emulators, which may not be affordable in computationally more demanding applications, such as three-body scattering (see Ref. [20, 21]). Second, it does not estimate errors.
@@ III.2.2 Emulation of phase shifts
Figure 4: High-fidelity (or FOM) and emulated (or ROM) phase shifts and mixing angles as a function of the laboratory energy for the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channel based on the GT+ potential. The panels (a) to (c) show the phase shifts in the $^{3}\text{S}_{1}$ (panel a) and in the $^{3}\text{D}_{1}$ channel (panel b), as well as the mixing angle $\varepsilon_{1}$ (panel c). The solid black lines represent the high-fidelity results, while the dots indicate the emulated results at a few selected energies. The panels (d) to (f) show the relative error (23) of the G-ROM and LSPG-ROM, respectively, with respect to the high-fidelity calculations. This comparison was made for two requested error tolerances on the $\tilde{t}-$matrix, $\alpha=10^{-4}$ and $10^{-7}$, as reported in the legend.
Next, we emulate phase shifts based on the GT+ potential. To this end, we construct emulators on a grid of laboratory energies for a given partial-wave channel. Figure 4 shows our results for the emulated phase shifts and mixing angles at the potential’s best-fit values of the LECs as a function of the laboratory energy in the coupled $^{3}\text{S}_{1}$$^{3}\text{D}_{1}$ channel. [^6] The best-fit values were obtained in Ref. [18] using a least-squares fit to scattering phase shifts and the $^{1}\text{S}_{0}$ scattering lengths. We choose the G-ROM and the deuteron channel as a coupled-channel case. The results are representative of the LSPG-ROM and other partial-wave channels we have studied, including the coupled and uncoupled P and F partial waves. [^7] Note that the approach in Ref. [28] was limited to uncoupled channels and local potentials (in coordinate space), restrictions that our emulators relax.
To obtain the emulated phase shifts in Fig. 4, we train the emulator using a pool of 500 candidate snapshots randomly selected in the region of $\pm 40\%$ of the best-fit LECs values obtained in Ref. [18]. Given a single initial snapshot randomly chosen from this pool, the greedy algorithm estimates the error of each remaining candidate snapshot, performs a single high-fidelity calculation for the one with the largest (estimated) error, and adds the orthonormalized snapshot to the emulator basis. This process is repeated until the requested error tolerance $\alpha$ specified in the legend is met. After training is complete, we run the emulators using the best-fit LEC values from Ref. [18].
@@
We conclude that the greedy algorithm enables efficient, accurate training of scattering emulators. Note that second-order corrections from applying the Kohn variational principle [15, 17] are not needed to achieve this level of accuracy.
In contrast to Ref. [28], we have not encountered spurious singularities known as Kohn (or Schwartz) anomalies that occur when the emulator equations are singular or near-singular [8]. This may suggest that scattering calculations in momentum space are less prone to such anomalies, regardless of the ROM used, likely because the $t$-matrix is complex-valued, as discussed in Ref. [8]. Only the LSPG-ROM has been analytically shown in Ref. [28] to be less prone to Kohn anomalies.